ASSOCIATED CURVES ON RULED SURFACES IN THE ISOTROPIC

SPACE I_{3}^{(1)} by Ivan Kamenarović, Rijeka, Croatia

Abstract. The paper deals with associated curves on ruled surfaces in the isotropic space I_{3}^{(1)}. On the base of previously found relations, the cases when these curves are lines of curvature or asymptotic lines are investigated. Eventually the case when one of these curves is a line of curvature and the other is an asymptotic line is treated too. Mathematics subject classification: 53A40. Keywords: Isotropic space, ruled surfaces, associated curves on ruled surfaces.